# Can a calibration metric be both testable and actionable?

### v1

@@ Abstract
Forecast probabilities often serve as critical inputs for binary decision making. In such settings, calibration—ensuring forecasted probabilities match empirical frequencies—is essential. Although the common notion of Expected Calibration Error (ECE) provides actionable insights for decision making, it is not testable: it cannot be empirically estimated in many practical cases. Conversely, the recently proposed Distance from Calibration (dCE) is testable but is not actionable since it lacks decision-theoretic guarantees needed for high-stakes applications. We introduce Cutoff Calibration Error, a calibration measure that bridges this gap by assessing calibration over intervals of forecasted probabilities. We show that Cutoff Calibration Error is both testable and actionable and examine its implications for popular post-hoc calibration methods, such as isotonic regression and Platt scaling.
@@ 1 Introduction
$\mathbb{E}[Y\mid f(X)]=f(X)\;\text{ almost surely.}$. (1)
@@
A challenge lies in quantifying how close the model $f$ comes to the ideal of perfect calibration (1). There are two main desiderata for any measure of approximate calibration $\Delta(f)$. First, it should be testable, in the sense that we should be able to reliably estimate the metric from finite data. In particular, we need to be able to test whether $f$ is approximately calibrated, $\Delta(f)\approx 0$. Second, any calibration metric should be actionable—if $\Delta(f)$ is small, this should imply meaningful guarantees for downstream decision-making. For example, if a medical diagnostic system’s probabilities have low calibration error, doctors should be confident that acting on its risk predictions according to clinical guidelines will lead to good patient outcomes. Unfortunately, as we will see, these two requirements are in tension: many natural measures of calibration that provide strong decision-theoretic guarantees turn out to be statistically intractable to estimate, and vice versa.
@@
$\Delta_{\operatorname{ECE}}(f):=\mathbb{E}\Big{|}\mathbb{E}\left[Y\middle|f(X)\right]-f(X)\Big{|}.$ (2)
@@
$\Delta_{\operatorname{dCE}}(f):=\inf_{\begin{subarray}{c}g:\mathcal{X}\rightarrow[0,1]\\ \mathbb{E}[Y\mid g(X)]=g(X)\end{subarray}}\mathbb{E}|g(X)-f(X)|.$ (3)
dCE is an easier notion of calibration to satisfy, as compared to ECE: indeed, for any $f$, $\Delta_{\operatorname{dCE}}(f)\leq\Delta_{\operatorname{ECE}}(f)$ (Błasiok et al., 2023, Lemma 4.7). Moreover, dCE solves the measurement challenge, since unlike ECE, it can be consistently estimated from data. However, this statistical advantage comes at a cost—dCE fails to provide the decision-theoretic assurances that made ECE valuable for applications. We show that models with small dCE can still lead to poor decisions.
@@
Figure 1: Comparing our proposal, Cutoff Calibration Error, to ECE and dCE.
To bridge the gap between testable and actionable calibration measures, we propose Cutoff Calibration Error, which assesses calibration over ranges of predicted probabilities rather than at specific values. We prove that Cutoff Calibration Error can be efficiently estimated from data at the parametric rate, while still providing guarantees about decision quality when using the calibrated forecasts. The key insight is that most practical decisions depend on whether a probability exceeds some threshold, making interval-based calibration particularly relevant.
@@
Popular existing measures face limitations (Sections 3 and 4).
We establish that testing ECE is fundamentally hard by establishing that any calibration method with asymptotically vanishing ECE must have small effective support size. Additionally, we demonstrate through counterexamples that dCE cannot provide the same decision-theoretic guarantees as ECE.
Cutoff Calibration Error is testable and actionable (Sections 2, 3 and 4):
We define Cutoff Calibration Error and demonstrate that it combines the strongest benefits of ECE and dCE. First, controlling Cutoff Calibration Error implies decision-theoretic guarantees in the binary decision setting under an intuitive monotonicity constraint. Second, like dCE, Cutoff Calibration Error can be efficiently estimated from data using a plug-in approach that’s closely related to a proposal of Arrieta-Ibarra et al. (2022). We study relationships between Cutoff Calibration Error, ECE and dCE in detail.
Implications for Platt scaling and isotonic regression (Section 5):
@@ 2 Our proposal: Cutoff calibration
Consider a medical diagnostic system that models probabilities of various conditions. While existing calibration measures like ECE focus on the accuracy of specific probability values (e.g., historical outcomes when the model output was exactly 73%), clinicians typically work with standardized risk thresholds—low, medium, and high risk categories that map to different intervention protocols. A doctor rarely needs to distinguish between a 73% and 74% risk, but they absolutely need to know if a patient falls above or below key decision thresholds, such as the 80% mark that might trigger immediate attention.
@@
$\Delta_{\operatorname{Cutoff}}(f):=\sup_{\textnormal{ intervals}\leavevmode\nobreak\ I}\Big{|}\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]\Big{|}.$ (4)
@@
How does Cutoff Calibration Error relate to the existing calibration metrics, ECE and dCE? We establish a precise first relationship through the following chain of inequalities.
@@
For any $f:[0,1]\rightarrow[0,1]$,
$\Delta_{\operatorname{dCE}}(f)^{2}/9\leq\Delta_{\operatorname{Cutoff}}(f)\leq\Delta_{\operatorname{ECE}}(f).$ (5)
@@ 2.1 Connections to weighted calibration error
$\Delta_{\operatorname{wCE}}(f;\mathcal{C}):=\sup_{w\in\mathcal{C}}\mathbb{E}[w(f(X))(Y-f(X))].$ (6)
@@
Figure 2: Relating our proposal, Cutoff Calibration Error, to ECE and dCE. Arrows—of the form $\Delta_{1}\Rightarrow\Delta_{2}$—signify that a calibration metric $\Delta_{2}$ can be bound in terms of the other $\Delta_{1}$.
@@
This form of weighted calibration error was previously known as weak calibration (Kakade and Foster, 2008) and smooth calibration (Gopalan et al., 2022). The connection to $1$-Lipschitz weighting functions lends further insight into why dCE may not be the most useful metric for decision making: since indicator functions are not Lipschitz, dCE is not compatible with hard thresholds. We further explore this issue in Section 3.
@@
Finally, note that $\Delta_{\operatorname{Cutoff}}(f)$ is already in the form of a weighted calibration error, where the weighting functions are indicator functions over intervals. Since the objective $\mathbb{E}[w(f(X))(Y-f(X))]$ in weighted calibration error is linear in $w$, we can replace this class with its convex hull, leading to a much richer class.
@@
Let $\mathcal{B}_{M}$ be the class of all functions on $[0,1]\rightarrow[-1,1]$ whose total variation is at most $M\in\mathbb{R}^{+}$. Then, for $M\geq 2$,
$\Delta_{\operatorname{Cutoff}}(f)\leq\Delta_{\operatorname{wCE}}(f;\mathcal{B}_{M})\leq 3\cdot(M+1)\Delta_{\operatorname{Cutoff}}(f).$ (7)
@@ 3.1 The simple binary decision setting
$\ell_{\mathrm{bd}}(Y,\widehat{Y};\tau)=\tau(1-Y)\widehat{Y}+(1-\tau)Y(1-\widehat{Y}),$ (8)
where $\widehat{Y}\in\{0,1\}$ encodes our preferred action if we knew $Y$.
The Bayes decision rule is ${{\mathbf{1}_{\left\{\mathbb{E}[Y\mid f(X)]\geq\tau\right\}}}}$. That is to say, the loss is minimized when the resources are deployed ($\widehat{Y}=1$) if and only if the calibrated forecast probability is above $\tau$. In practice, $\mathbb{E}[Y\mid f(X)]$ is unknown, so decision-makers naturally resort to the following plug-in decision rule:
@@
The question remains how good it is to use $f(X)$ as a proxy for $\mathbb{E}[Y\mid f(X)]$. We measure this via the risk (expected loss): define $R_{\mathrm{bd}}:\mathcal{F}\times[0,1]\rightarrow\mathbb{R}$ to be the risk (expected loss) of a forecast function $f\in\mathcal{F}$ and parameter $\tau\in[0,1]$:
@@ 3.2 Comparing calibration measures through decision guarantees
Suppose $\tau=0.75$ and we only have access to $\tilde{f}(X)$. Then, the Bayes decision rule will be to use $\widehat{Y}={{\mathbf{1}_{\left\{\mathbb{E}[Y\mid\tilde{f}(X)]\geq 0.75\right\}}}}$. If we use ${{\mathbf{1}_{\left\{\tilde{f}(X)\geq 0.75\right\}}}}$ in place of the Bayes decision rule, then the risk $R_{\mathrm{bd}}(\tilde{f}(X);\tau)]=\frac{3}{8}$. However, the Bayes decision rule actually has 0 risk, since in fact $Y={{\mathbf{1}_{\left\{\mathbb{E}[Y\mid\tilde{f}(X)]\geq 0.75\right\}}}}$. Therefore, while $\Delta_{\operatorname{dCE}}(\tilde{f})$ can be arbitrarily small, the discrepancy in losses between the plug-in decision rule and the oracular Bayes decision rule remains bounded away from 0.
When ECE is small, we can guarantee that the risk associated with $f$ is close to the best possible wrapper $h\circ f$. We often refer to this difference as the “risk gap.”
@@
$R_{\mathrm{bd}}(f;\tau)-\inf_{h:[0,1]\rightarrow[0,1]}R_{\mathrm{bd}}(h\circ f;\tau)\leq\Delta_{\operatorname{ECE}}(f).$ (10)
@@
Hu and Wu (2024) examined this problem more generally, where instead of considering our specific loss function they worked with all bounded loss functions that are proper scoring rules. They propose a measure of calibration that bounds this risk gap (often termed “swap regret” in game theory). Their proposed definition of calibration ends up resembling the right hand side of (11), with a supremum taken over $h$ and $\tau$, using steps found also in Kleinberg et al. (2023); Li et al. (2022). While their proposed measure of calibration is $\leq\Delta_{\operatorname{ECE}}(f)$, it is also $\geq\Delta_{\operatorname{ECE}}(f)^{2}$ (Hu and Wu, 2024, Theorem 4.1). Unfortunately, in the continuous random variable $f(X)$ context we focus on in our work, this means that the proposed calibration measure of Hu and Wu (2024) is not testable, like ECE.
@@
A natural shape constraint on $h$ that comports with common sense is that it be monotonic. By constraining ourselves to monotonic wrappers around $f$, we can upper bound the monotone risk gap by Cutoff Calibration Error.
@@
$R_{\mathrm{bd}}(f;\tau)-\inf_{\begin{subarray}{c}h:[0,1]\rightarrow[0,1]\\ \textnormal{monotone}\end{subarray}}R_{\mathrm{bd}}(h\circ f;\tau)\leq 2\Delta_{\operatorname{Cutoff}}(f).$ (12)
While we state this result just for binary decision loss, a corollary is that we can bound the risk gaps for arbitrary bounded proper scoring rules when $Y\in\{0,1\}$ (Appendix B). This corollary follows from the Schervish representation (Schervish, 1989) of proper scoring rules, which demonstrates that proper scoring rules are a mixture of binary decision loss across different $\tau$ values.
@@ 4 Testable calibration: Distribution-free guarantees
In order for a calibration metric to be useful for decision-making in practice, as discussed in Section 3, we need to choose a metric $\Delta$ that is actionable. But this is not sufficient: we also need to be able to verify empirically that $\Delta^{P}(f_{n})\approx 0$, [^2] When we write $\Delta^{P}(f_{n})$, we proceed as if $f_{n}$ were fixed. We use the $P$ superscript to emphasize that the expectation is only taken over $P$, not $P^{n}$. E.g., $\Delta_{\operatorname{wCE}}^{P}(f_{n})=\sup_{w\in\mathcal{C}}\mathbb{E}_{P}[(Y-f_{n}(X)){{\mathbf{1}_{\left\{f_{n}(X)\right\}}}}\mid f_{n}]$. in order to know whether our particular trained model $f_{n}$ can then be safely used for decision-making. In other words, for a small constant $c>0$,
@@
Of course, this is always possible with a trivial solution: we can simply return a constant function $f_{n,\textnormal{const}}(x)\equiv\hat{\mu}_{Y}$ where $\hat{\mu}_{Y}$ is the sample mean of $Y$ (since we can always estimate the mean $\mathbb{E}Y$ with error $\mathcal{O}_{P}(n^{-1/2})$, this will ensure that $\Delta^{P}(f_{n,\textnormal{const}})$ is low). We would ideally want to use a state-of-the-art fitted model $f_{n}$, not some overly simple model such as a constant function. We can therefore ask a related question:

### v2

@@ Abstract
Forecast probabilities often serve as critical inputs for binary decision making. In such settings, calibration—ensuring forecasted probabilities match empirical frequencies—is essential. Although the common notion of Expected Calibration Error (ECE) provides actionable insights for decision making, it is not testable: it cannot be empirically estimated in many practical cases. Conversely, the recently proposed Distance from Calibration (dCE) is testable, but it is not actionable since it lacks decision-theoretic guarantees needed for high-stakes applications. To resolve this question, we consider Cutoff Calibration Error, a calibration measure that bridges this gap by assessing calibration over intervals of forecasted probabilities. We show that Cutoff Calibration Error is both testable and actionable, and we examine its implications for popular post-hoc calibration methods, such as isotonic regression and Platt scaling.
@@ 1 Introduction
$\mathbb{E}[Y\mid f(X)]=f(X)\;\text{ almost surely.}$ (1)
@@
A challenge lies in quantifying how close the model $f$ comes to the ideal of perfect calibration (1). There are two main desiderata for any measure of approximate calibration $\Delta(f)$. First, it should be testable, in the sense that we should be able to reliably estimate the metric from finite data. In particular, we need to be able to test whether $f$ is approximately calibrated, $\Delta(f)\approx 0$. Second, any calibration metric should be actionable—if $\Delta(f)$ is small, this should imply meaningful guarantees for downstream decision-making. For example, if a medical diagnostic system’s probabilities have low calibration error, doctors should be confident that acting on its risk predictions according to clinical guidelines will lead to good patient outcomes. Unfortunately, as we will see, these two requirements are in tension: natural measures of calibration that provide strong decision-theoretic guarantees often turn out to be statistically intractable to estimate, and vice versa. We discuss in more detail these desiderata of actionability and testability in Section 3 and Section 4, respectively.
@@
$\Delta_{\operatorname{ECE}}(f):=\mathbb{E}\left[\Big{|}\mathbb{E}\left[Y\middle|f(X)\right]-f(X)\Big{|}\right].$
@@
$\Delta_{\operatorname{dCE}}(f):=\inf_{\begin{subarray}{c}g:\mathcal{X}\rightarrow[0,1]\\ \mathbb{E}[Y\mid g(X)]=g(X)\end{subarray}}\mathbb{E}\left[|g(X)-f(X)|\right].$
dCE is an easier notion of calibration to satisfy, as compared to ECE: indeed, for any $f$, $\Delta_{\operatorname{dCE}}(f)\leq\Delta_{\operatorname{ECE}}(f)$ (Błasiok et al., 2023, Lemma 4.7). Moreover, dCE solves the issue of testability, since unlike ECE, bounds on it can be estimated from data. However, this statistical advantage comes at a cost—dCE fails to provide the decision-theoretic assurances that made ECE valuable for applications. We show that models with small dCE can still lead to poor decisions.
@@
Figure 1: Comparing Cutoff Calibration Error to ECE and dCE.
To bridge the gap between testable and actionable calibration measures, we consider the Cutoff Calibration Error, which we can view as an intermediate definition lying between dCE and ECE. This measure assesses calibration over ranges of predicted probabilities rather than at specific values. We prove that Cutoff Calibration Error can be efficiently estimated from data at the parametric rate, while still providing guarantees about decision quality when using the calibrated forecasts. The key insight is that most practical decisions depend on whether a probability exceeds some threshold, making interval-based calibration particularly relevant.
@@
Popular existing measures face limitations.
Building on the results of Gupta et al. (2020), we establish that testing ECE is fundamentally hard by establishing that any calibration method with asymptotically vanishing ECE must have small effective support size. On the other hand, we demonstrate through counterexamples that dCE cannot provide the same decision-theoretic guarantees as ECE. In other words, we will see that ECE is not testable, while dCE is not actionable.
Cutoff Calibration Error is testable and actionable:
We define Cutoff Calibration Error and demonstrate that it combines the strongest benefits of ECE and dCE. First, controlling Cutoff Calibration Error implies decision-theoretic guarantees in the binary decision setting under an intuitive monotonicity constraint. Second, like dCE, Cutoff Calibration Error can be efficiently estimated from data using a plug-in approach that’s closely related to a proposal of Arrieta-Ibarra et al. (2022). We study relationships between Cutoff Calibration Error, ECE and dCE in detail. [^1] After submission of this paper, we learned of two concurrent works (Okoroafor et al., 2025; Qiao and Zhao, 2025) that address similar questions. Each proposes related definitions of calibration, and each considers the benefits of these definitions from the perspectives of a decision-theoretic framework, albeit in distinct ways; we will comment on these similarities and distinctions as appropriate when presenting our results below. In addition, Okoroafor et al. (2025) also considers testability, although our paper focuses more on how testability hardness results motivate the construction of Cutoff Calibration Error.
Implications for Platt scaling and isotonic regression:
@@ 2 Our proposal: Cutoff calibration
Consider a medical diagnostic system that models probabilities of various conditions. While existing calibration measures like ECE focus on the accuracy of specific probability values (e.g., historical outcomes when the model output was exactly 73%), clinicians typically work with standardized risk thresholds—low, medium, and high risk categories that map to different intervention protocols. A doctor rarely needs to distinguish between a 73% and 74% risk, but they need to know whether a patient falls above or below key decision thresholds; perhaps the 80% mark may prompt immediate attention.
@@
$\Delta_{\operatorname{Cutoff}}(f):=\sup_{\textnormal{ intervals}\leavevmode\nobreak\ I}\Big{|}\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]\Big{|}.$
@@
How does Cutoff Calibration Error relate to the existing calibration metrics, ECE and dCE? We establish a precise relationship through the following chain of inequalities.
@@
For any $f:[0,1]\to[0,1]$,
$\Delta_{\operatorname{dCE}}(f)^{2}/9\leq\Delta_{\operatorname{Cutoff}}(f)\leq\Delta_{\operatorname{ECE}}(f).$
@@ 2.1 Connections to weighted calibration error
$\Delta_{\operatorname{wCE}}(f;\mathcal{C}):=\sup_{w\in\mathcal{C}}\mathbb{E}[w(f(X))(Y-f(X))].$
@@
Figure 2: Relating Cutoff Calibration Error to ECE and dCE. Arrows—of the form $\Delta_{1}\Rightarrow\Delta_{2}$—signify that a calibration metric $\Delta_{2}$ can be bound in terms of $\Delta_{1}$.
@@
This form of weighted calibration error was previously known as weak calibration (Kakade and Foster, 2008) and smooth calibration (Gopalan et al., 2022). The connection to $1$-Lipschitz weighting functions lends further insight into why dCE may not be the most useful metric for decision making: since indicator functions are not Lipschitz, dCE is not compatible with hard thresholding rules. We further explore this issue in Section 3.
@@
Finally, note that $\Delta_{\operatorname{Cutoff}}(f)$ is already in the form of a weighted calibration error, where the weighting functions are indicator functions over intervals. Since the objective
$\mathbb{E}[w(f(X))(Y-f(X))]$
in weighted calibration error is linear in $w$, we can replace this class with its convex hull, leading to a much richer class.
@@
Let $\mathcal{B}_{M}$ be the class of all functions on $0,1]\rightarrow[-1,1]$ whose total variation is at most $M\in\mathbb{R}^{+}$. Then, for $M\geq 2$,
$\Delta_{\operatorname{Cutoff}}(f)\leq\Delta_{\operatorname{wCE}}(f;\mathcal{B}_{M})\leq(M+2)\Delta_{\operatorname{Cutoff}}(f).$
@@ 3.1 The simple binary decision setting
[…]
Given an observed forecast $f(X)$, the Bayes decision rule is given by $\mathbf{1}_{\left\{\mathbb{E}[Y\mid f(X)]\geq\tau\right\}}$}}. That is to say, the expected loss is minimized when the resources are deployed ($\widehat{Y}=1$) if and only if the calibrated forecast probability is above $\tau$. In practice, $\mathbb{E}[Y\mid f(X)]$ is unknown, so decision-makers naturally resort to the following plug-in decision rule:
@@
But, if $f$ is highly miscalibrated, this might be a poor choice of decision rule.
Examining this example, we can see that we are asking whether the decision rule $\mathbf{1}_{\left\{f(X)\geq\tau\right\}}$}} is (nearly) as good as some other decision rule obtained by transforming $f$, $\mathbf{1}_{\left\{h(f(X))\geq\tau\right\}}$}} (observe that the Bayes rule can be written in this form by taking $h(t)=\mathbb{E}[Y\mid f(X)=t]$). We are particularly interested in whether this favorable property is ensured for functions $f$ that satisfy $\Delta(f)\approx 0$, given some particular choice of calibration measure $\Delta$. This is what we mean by saying that $\Delta$ is actionable: we are asking whether low miscalibration ($\Delta(f)\approx 0$) implies that the plug-in decision rule is nearly optimal relative to some class of possible modifications of the rule—that is, some set of transformations $\mathcal{H}$.
@@ 3.2 Comparing calibration measures through decision guarantees
Before presenting the example, we first define a measure of risk for the plug-in decision rule:
@@
[…]
From this example, we can see that a small dCE error, $\Delta_{\operatorname{dCE}}(f)\approx 0$, does not necessarily ensure that the forecast $f$ is actionable. [^2] Interestingly, there is a certain sense in which dCE is “close” to being actionable: adding a small amount of noise to a forecast $f(X)$ with small dCE will lead to an actionable rule. In fact, though, this is because $f(X)+\mathrm{noise}$ is then guaranteed to have small ECE (Blasiok and Nakkiran, 2024). In other words, it is possible to build a low-ECE (and therefore, actionable) forecast by using the low-dCE rule $f(X)$ as a starting point—but $f(X)$ is not itself actionable, and decisions made by thresholding $f(X)+\mathrm{noise}$ might be arbitrarily different from decisions made using $f(X)$.
In contrast, when ECE is small, we can directly guarantee that the risk associated with $f$ is close to the best possible wrapper $h\circ f$. We often refer to this difference as the “risk gap.”
@@
$R_{\mathrm{bd}}(f;\tau)-\inf_{h:[0,1]\rightarrow[0,1]}R_{\mathrm{bd}}(h\circ f;\tau)\leq\Delta_{\operatorname{ECE}}(f).$
@@
Proposition 3.1 also follows from a more general result in Hu and Wu (2024), where instead of considering our specific loss function they worked with all bounded loss functions that are proper scoring rules. They propose a measure of calibration that bounds this risk gap (often termed “swap regret” in game theory). Their proposed definition of calibration ends up resembling the right hand side of (3), with a supremum taken over $h$ and $\tau$, using steps found also in Kleinberg et al. (2023); Li et al. (2022). While their proposed measure of calibration is $\leq\Delta_{\operatorname{ECE}}(f)$, it is also $\geq\Delta_{\operatorname{ECE}}(f)^{2}$ (Hu and Wu, 2024, Theorem 4.1). Unfortunately, in the setting of a continuous random variable $f(X)$ that is the focus of our work, this means that the proposed calibration measure of Hu and Wu (2024) is not testable, like ECE.
In parallel, Kleinberg et al. (2023) leverage similar observations and propose a calibration metric that is in fact testable. However, its guarantee is for external regret, which compares against a non-personalized treatment policy. Thus, the metric proposed in Kleinberg et al. (2023) is not actionable in the sense that we define here in our work. Relatedly, Qiao and Zhao (2025) consider decision-theoretic guarantees from a different perspective, including benchmarking against non-personalized treatment policies.
@@
A natural shape constraint on $h$ that comports with common sense is that it be monotonically non-decreasing. By constraining ourselves to monotonic wrappers around $f$, we can upper bound the monotone risk gap by Cutoff Calibration Error.
@@
$R_{\mathrm{bd}}(f;\tau)-\inf_{\begin{subarray}{c}h:[0,1]\rightarrow[0,1]\\ \textnormal{monotone}\end{subarray}}R_{\mathrm{bd}}(h\circ f;\tau)\leq 2\Delta_{\operatorname{Cutoff}}(f).$
An alternate way of proving this result, up to constant factors, is to use the results of Okoroafor et al. (2025, Lemmas 3.2 and 3.5), who examine this question from the perspective of omniprediction (Gopalan et al., 2021). The salient fact in both our proof and one leveraging results in Okoroafor et al. (2025) is that pre-images of super-level sets of univariate monotone functions are half-spaces.
While we state Proposition 3.2 just for binary decision loss, a corollary is that we can bound the risk gaps for arbitrary bounded proper scoring rules when $Y\in\{0,1\}$ (Appendix B). This corollary follows from the Schervish representation (Schervish, 1989) of proper scoring rules, which demonstrates that proper scoring rules are a mixture of binary decision loss across different $\tau$ values.
@@ 4 Testable calibration: Distribution-free guarantees
In order for a calibration metric to be useful for decision-making in practice, as discussed in Section 3, we need to choose a metric $\Delta$ that is actionable. But this is not sufficient: we also need to be able to verify empirically that $\Delta^{P}(f_{n})\approx 0$, [^4] When we write $\Delta^{P}(f_{n})$, we proceed as if $f_{n}$ were fixed. We use the $P$ superscript to emphasize that the expectation is only taken over $P$, not over the data sampled from $P^{n}$ and used to train $f_{n}$ (that is, the expectation is computed conditional on $f_{n}$). In particular for a fixed function $f$, $\Delta(f)=\Delta^{P}(f)$, since conditioning on the fixed function $f$ would have no effect. in order to know whether our particular trained model $f_{n}$ can then be safely used for decision-making. In other words, for a small constant $c>0$,
@@
Of course, this is always possible with a trivial solution: we can simply return a constant function $f_{n,\textnormal{const}}(x)\equiv\hat{\mu}_{Y}$ where $\hat{\mu}_{Y}$ is the sample mean of $Y$ (since we can always estimate the mean $\mathbb{E}[Y]$ with error $\mathcal{O}_{P}(n^{-1/2})$, this will ensure that $\Delta^{P}(f_{n,\textnormal{const}})$ is low). However, we would ideally want to use a state-of-the-art fitted model $f_{n}$, not some overly simple model such as a constant function. We can therefore ask a related question: